If a spring is doubled in length, then the potential energy stored in the spring will be
1
Expert's answer
2020-12-14T07:25:07-0500

Let`s solve this question:

we need to define K if length of spring is doubled:

"k=E\times \frac{S}{l}", E - young modules, S- area of spring and l-length of spring, So if we doubled length:

"k1=E\times \frac{S}{2\times l}"="\frac{k}{2}", k decreases twice than first vaue.

Potential energy: "E=\frac{k1\times l^2}{2}" if we put k and l:

"E(final)=\frac{\frac{k}{2}\times (2\times l)^2}{2}=2\times E" this means if doubled length of spring, then its potential energy also doubled.